### v1

@@ 3 Methods
I present a self-consistent framework for PopSynth emulation equipped with well-calibrated uncertainty quantification and model comparison, which can be used for reliable astrophysical inference and feature production using state-of-the-art simulators and growing catalogs of observations. PopSynth simulations generate an ensemble of systems with uniquely distributed parameters ($\vec{\theta}$) given a particular set of initial conditions and/or other quantities ($\vec{\lambda}$) that characterize the evolutionary track being simulated. Hence, emulating PopSynth essentially amounts to constructing accurate estimators for the conditional distribution function:
@@ 3.1 Normalizing Flows
Here $\vec{\omega}$ are the parameters of the DNN, $p_{u}$ is a base distribution comprised of i.i.d. standard Gaussian, and $f_{NN}$ is a sequence of differentiable and invertible transformations. Once the model is trained to find an optimal $\vec{\omega}=\vec{\omega}_{0}$, samples can be drawn from an emulated population in between the training point $(\vec{\lambda}=\vec{\lambda}_{new})$ by first drawing $\vec{u}_{i}\sim p_{u}$ and then computing $\vec{\theta}_{i}=f^{-1}_{NN}(\vec{u}_{i},\vec{\lambda}_{new},\vec{\omega}_{0})$. Data amplification for training points is also possible for boosting sample-statistics in under-represented subpopulations for improving feature prediction. Similarly, the conditional density becomes tractable as a numerical function suitable for repeated evaluations, also in between training points $(\vec{\theta}=\vec{\theta}_{new},\vec{\lambda}=\vec{\lambda}_{new})$, which is required to implement SBI from observed catalogs (S. Colloms et al., 2025).
Various architectures have been proposed in literature for constructing $f_{NN}$, each with its own merits and demerits. Broadly, the transformation functions can be classified into two categories: discrete layers of finite depth, such as masked autoregressive flows (MAF, G. Papamakarios et al., 2018), and continuous time transforms, which are essentially diffusion models that asymptotically approach the target density (C. Chen et al., 2018; W. Grathwohl et al., 2018). While various sub-categories under these two classes can be advantageous to use for specific problems, continuous normalizing flow (CNF) algorithms such as FFJORD (W. Grathwohl et al., 2018) are, in general, more scalable and require DNNs of much lower complexity to obtain the same expressivity as a discrete flow like MAF. I shall show this in the context of our PopSynth datasets while noting that the lower complexity makes CNFs ideal for the Bayesian framework proposed for UQ.
For any given transformation architecture, I can train the NF model by minimizing some loss function with respect to $\vec{\omega}$, such that the optimal $\vec{\omega}$ corresponds to $p$ and $\hat{p}$ being maximally similar (D. J. Rezende & S. Mohamed, 2016). As a standard metric, the expected KL divergence (averaged over the list of simulations on which I want to train) is chosen. Assuming the training simulations where carried out for gridpoints: $\{(\vec{\lambda})\sim p_{\lambda}(\vec{\lambda})\}$, the training dataset: $\vec{d}_{train}=\{(\vec{\theta}_{i},\vec{\lambda}_{i})\sim p(\vec{\theta},\vec{\lambda})\}_{i=1}^{N}$ can be thought of as a set of i.i.d. samples drawn from the joint distribution $p(\vec{\theta},\vec{\lambda})=p(\vec{\theta}|\vec{\lambda})p_{\lambda}(\vec{\lambda})$. Under these considerations, the expected KL divergence can be shown to equal the following sum over training datapoints (up to an additive constant):
@@
which can be used for tractable density evaluation, data generation, and amplification. In other words, the trained emulator can, in principle, facilitate SBI and feature prediction for state-of-the-art simulators such as POSYDONv2.
The KL divergence as a function of $\vec{\omega}$ can be complicated with multiple peaks, troughs, and saddle points, leading to gradient-based minimization algorithms finding slightly different answers for the best fit estimator every time the model is re-trained. Loss function morphology can also change significantly between alternate realizations of thinned and noisy training data. In other words, if I train the same model on the same dataset twice, or on different realizations of a noisy dataset, I won’t necessarily get the exact same answer for the predicted population. It is important to quantify and marginalize over this variance in flow predictions for accurate SBI and feature prediction, particularly in the cases of sparse and noisy training datasets and sub-optimal architecture choices. Hence, instead of a best-fit emulator that deterministically yields a single distribution, it is desirable to predict a range of density functions that are likely to contain the true population with a certain posterior probability given training data, model assumptions, and a priori expectation on DNN parameters. Before I show how that can be achieved, it is worthwhile to discuss the types and sources of these uncertainties that I wish to quantify in a bit more detail.
@@ 3.2 Uncertainties in flow predictions
On the other hand, uncertainty can also arise from the inherent randomness, sparsity, and class-confusion present in the dataset, which is irreducible even with the addition of more training samples (W. He et al., 2025; V. Nemani et al., 2023; E. Hüllermeier & W. Waegeman, 2021). Two key examples of these aleatoric/data uncertainties in the PopSynth of merging compact binaries are the sparsity of certain regions of parameter space, such as extreme-mass ratio binaries with “mass-gap” compact objects (M. Zevin et al., 2020; R. Abbott et al., 2020) where the merger rate is expected to be low, and degenerate population features shared by different values of $\vec{\lambda}$. The former would become particularly relevant if the methods of S. Colloms et al. (2025) were to be applied to mass-gap NSBH-like systems, of which two have already been observed (R. Abbott et al., 2020; A. G. Abac et al., 2024). Quantifying data uncertainties in deterministic DNN predictors can be achieved through Bayesian inference (J. Arbel et al., 2023) and parametric or non-parametric modeling of the distribution of noise in the data, as well as the use of deep probabilistic models such as secondary NFs and generative adversarial networks (E. Hüllermeier & W. Waegeman, 2021; J. Whang et al., 2021; S. Dirmeier et al., 2023). Again, such approaches have not been popularly used to quantify the uncertainties of NF parameters, which is relevant in scenarios wherein the NF itself is the underlying DNN predictor (see L. Berry & D. Meger, 2023, , for an implementation of Monte Carlo dropout based ensemble network approach on NFs).
In the following, I introduce the Bayesian NF, a generative probabilistic model constructed from BNNs, which can self-consistently quantify epistemic uncertainties that arise from inefficient training and aleatoric uncertainties that arise from the sparsity and stochasticity inherent to the dataset. No approximations for the log-likelihood function are utilized, and rigorous sampling methods are used for implementing Bayesian inference. The resulting predictions for an emulated PopSynth simulation are hence marginalized over such uncertainties and also automatically enable Bayesian comparison among different NF models, thereby minimizing the epistemic uncertainties that arise from choosing sub-optimal architectures and families of networks.
3.3 Uncertainty Quantification and Model Comparison: Bayesian Normalizing Flows
To quantify these epistemic and aleatoric uncertainties in the NF model for thinned datasets, I propose the following. As can be seen from Eqn. (3), The KL divergence is proportional to the negative log-likelihood of the model parameters given the training dataset (up to an additive constant):
@@ 3.3 Uncertainty Quantification and Model Comparison: Bayesian Normalizing Flows
where $\log\hat{p}(\vec{\theta},\vec{\lambda}|\vec{\omega})=\log\hat{p}(\vec{\theta}|\vec{\lambda},\vec{\omega})+\log p_{\lambda}(\vec{\lambda})$. Note that this is the exact log-likelihood of the parameters characterizing a probabilistic (NF) model and no a priori assumption about the distribution of noise in the dataset is necessary to construct its functional form. The training method in Sec. 3.1 above, therefore, amounts to finding a maximum likelihood density estimate (MLE). To quantify uncertainties, I instead propose to sample the posterior distribution of the model parameters using some suitably chosen prior and the exact log-likelihood:
@@
where m stands for marginalized. This amounts to constructing and training a normalizing flow whose transformations comprise a BNN. The predicted population will now represent a credible interval instead of a single distribution, which is expected to contain the true population with a certain posterior probability, given training data and prior expectations on the model parameters.
@@ 3.4 Prior choices, Implementation and Scalability
Choosing the prior distribution on model parameters correctly is crucial for convergent and scalable Bayesian inference, particularly for models of such high complexity. To fully explore the likelihood function, highly restrictive priors should be avoided. However, broad priors such as standard Gaussians or broad uniform distributions might lead to slow convergence for highly complex probabilistic models such as discrete NFs that are usually characterized by $O(10^{5}-10^{6})$ parameters. For highly complex models such as MAFs, it is instead desirable to choose a prior that is only supportive near the MLE parameters and yet broad enough for likelihood-driven inference. Furthermore, instead of fixing the width of the prior, it can also be allowed it to vary within a certain range. I chose the following simple choice for the prior distribution for this proof-of-concept demonstration.
@@
I start by choosing a maximal scale $\vec{\sigma}=\{\sigma_{0}\}_{i=1}^{n_{\omega}}$, which amounts to assuming a priori that the highest likelihood model parameters are within a factor of at most $\sigma_{0}$ of $\vec{\omega}_{0}$. I then define the standardized model parameters $\vec{\omega}_{st}\underset{i.i.d}{\sim}\mathcal{U}(-1,1)$ which I assume to be distributed uniformly within $(-1,1)$. I then rescale these to become numbers that lie within a factor of $\vec{\sigma}$ of $\vec{\omega}_{0}$ and hence construct the prior on $\vec{\omega}=\vec{\omega}_{0}(1+\vec{\sigma}\vec{\omega}_{st})$ as an affine transformed uniform distribution. The prior scale $\vec{\sigma}$ can, in principle, be sampled as well from a uniform or half-normal distribution, but for simplicity, I fix it to a value that allows for a reasonably broad prior and likelihood-driven inference. As I demonstrate shortly, calibration statistics can be constructed for choosing optimal values of $\sigma_{0}$.
@@
While these priors might seem restrictive as compared to a standard Gaussian, they can be expected to suffice as long as the posterior is found to be tightly constrained far away from the prior boundaries. I show that this is the case in the context of our population synthesis dataset and suggest increasing $\sigma_{0}$ for scenarios wherein the posterior is found to be railing against the prior. More sophisticated prior choices that are less restrictive and yet consistent with fast convergence are left as a future exploration.
@@
Even with these prior choices, the Bayesian NFs can quickly become intractable for high-dimensional distributions since discrete flows like MAF require highly complex networks to accurately capture all the features in the population. Bayesian inference can then amount to exploring (and storing) potentially thousands of copies of a million-parameter model, which is expected to become prohibitive. As I show later on in appendix A.1, CNFs can solve this problem by reducing the required model complexity for accurate density estimation by several orders of magnitude. Since the proposed Bayesian framework is agnostic of model architecture, it can be expected to work with CNFs in the same way as discrete flows, and thereby quantify uncertainties in the emulators of higher-dimensional population distributions. On the other hand, if reducing model complexity through the use of CNFs is not a viable alternative, or if complex embedding nets are necessary to reduce the dimensionality of $\vec{\lambda}$, approximate Bayesian inference can be implemented through a combination of stochastic variational inference and importance sampling which is outlined in appendix A.2.
I leave the full realization of both of these resolutions for the future and restrict to a Bayesian MAF that emulates a two-dimensional synthetic population in this proof-of-concept demonstration.
3.5 Calibration
@@
To assess the accuracy of the estimated uncertainties and optimize the choice of $\sigma_{0}$, I construct metrics that compare the empirical coverage of a test set to the theoretical coverage predicted by Bayesian inference (S. Bieringer et al., 2024). After drawing M posterior samples of flow parameters I split them up into m random subsets: $S_{\vec{\omega}}=\bigcup_{i=1}^{m}S_{\vec{\omega}}^{i}$, each representative of a smaller Bayesian inference run. For every posterior sample in a particular subset, I generate a million binaries $\vec{\theta}\in S_{\vec{\theta}}^{ij}=\{\vec{\theta}_{k}\sim\hat{p}(\vec{\theta}|\vec{\lambda}_{test},\vec{\omega}_{j})|\vec{\omega}_{j}\in S_{\vec{\omega}}^{i}\}_{k=1}^{10^{6}}$. I then construct $n_{Q}$ equal probability mass $\vec{\theta}$ bins (quantile binning) from the test set and compute the count of predicted samples in each bin. I then compute credible intervals for the counts and estimate the empirical coverage as the fraction of times (out of m and averaged over bins) the credible interval encloses the actual count of test samples. The calibration curve is then constructed as the relationship between the empirical coverage and the nominal coverage (confidence level of the credible intervals, S. Bieringer et al., 2024). Note, however, that for very large values of $\sigma_{0}$, if the walkers get stuck at a local maxima of the likelihood function, there will be biases resulting in low coverage. In other words, low coverage can be indicative of both under-estimation of uncertainties due to small $\sigma_{0}$ or biases due to very large $\sigma_{0}$ causing difficulty in convergence. This issue was not addressed by (S. Bieringer et al., 2024)
@@
1:$(N,d_{\theta},d_{\lambda})$: data dimensions; $\vec{d}_{train}\in\mathbb{R}^{N\times(d_{\theta}+d_{\lambda})}$: training data
$(\textsc{Adam},\alpha,\beta_{1},\beta_{2})$: Optimizer and associated parameters $\triangleright$ From D. P. Kingma & J. Ba (2017)
3:$(\textsc{NUTS},\epsilon,M)$: Sampler and associated parameters $\triangleright$ From M. D. Hoffman & A. Gelman (2011)
4:$F(\vec{\theta},\vec{\lambda},\vec{\omega})$: Flow model, computes $\hat{p}(\vec{\theta}|\vec{\lambda},\vec{\omega})$
5:$n_{\omega}$: model complexity, $\vec{\sigma}$: fixed prior scale for each model parameter
6:$avg\in[\mathrm{True},\mathrm{False}]$: whether to use the average or exact log-likelihood
7:$S_{\vec{\omega}},\text{BIC}$
8:$\vec{\omega}_{init}\sim\mathcal{N}(0,I)$; $\vec{\omega}_{0}\leftarrow\textsc{Adam}(\alpha,\beta_{1},\beta_{2},\textsc{Loss},\vec{\omega}_{init})$
9:$S_{\vec{\omega}}\leftarrow\textsc{NUTS}(\vec{\omega}_{init},\epsilon,\textsc{LogPost},M)$
10:$L_{max}\leftarrow-\infty$
@@
20:     return $P$if $\frac{\vec{\omega}-\vec{\omega}_{0}}{\vec{\sigma}\times\vec{\omega}_{0}(\textsc{elementwise})}\in[-1,1]$else $\infty$- ∞
@@
23:     return $-\textsc{LogLikelihood}(\vec{\omega},\mathrm{True})$

### v2

@@ 5 Normalizing Flow Emulators for population synthesis simulations
We start by noting that PopSynth simulations generate an ensemble of systems with uniquely distributed parameters ($\vec{\theta}$) given a particular set of initial conditions and/or other quantities ($\vec{\lambda}$) that characterize the evolutionary track being simulated. Hence, emulating PopSynth essentially amounts to constructing accurate estimators for the conditional distribution function:
@@
Here $\vec{\omega}$ are the parameters of the DNN, $p_{u}$ is a base distribution comprised of i.i.d. standard Gaussian, and $f_{NN}$ is a sequence of differentiable and invertible transformations. Once the model is trained to find an optimal $\vec{\omega}=\vec{\omega}_{0}$, samples can be drawn from an emulated population in between the training point $(\vec{\lambda}=\vec{\lambda}_{new})$ by first drawing $\vec{u}_{i}\sim p_{u}$ and then computing $\vec{\theta}_{i}=f^{-1}_{NN}(\vec{u}_{i},\vec{\lambda}_{new},\vec{\omega}_{0})$. Data augmentation for training points is also possible for boosting sample-statistics in under-represented subpopulations for improving feature prediction. Similarly, the conditional density becomes tractable as a numerical function suitable for repeated evaluations, also in between training points $(\vec{\theta}=\vec{\theta}_{new},\vec{\lambda}=\vec{\lambda}_{new})$, which is required to implement SBI from observed catalogs (S. Colloms et al., 2025).
Various architectures have been proposed in literature for constructing $f_{NN}$, each with its own merits and demerits. Broadly, the transformation functions can be classified into two categories: discrete layers of finite depth, such as masked autoregressive flows (MAF, G. Papamakarios et al., 2018), and continuous time transforms, which are essentially diffusion models that asymptotically approach the target density (C. Chen et al., 2018; W. Grathwohl et al., 2018). While various sub-categories under these two classes can be advantageous to use for specific problems, continuous normalizing flow (CNF) algorithms such as FFJORD (W. Grathwohl et al., 2018) are, in general, more scalable and require DNNs of much lower complexity to obtain the same expressivity as a discrete flow like MAF. We shall show this in the context of our PopSynth datasets while noting that the lower complexity makes CNFs ideal for the Bayesian framework proposed for UQ.
For any given transformation architecture, we can train the NF model by minimizing some loss function with respect to $\vec{\omega}$, such that the optimal $\vec{\omega}$ corresponds to $p$ and $\hat{p}$ being maximally similar (D. J. Rezende & S. Mohamed, 2016). As a standard metric, the expected KL divergence (averaged over the list of simulations on which we want to train) is chosen. Assuming the training simulations where carried out for gridpoints: $\{(\vec{\lambda})\sim p_{\lambda}(\vec{\lambda})\}$, the training dataset: $\vec{d}_{train}=\{(\vec{\theta}_{i},\vec{\lambda}_{i})\sim p(\vec{\theta},\vec{\lambda})\}_{i=1}^{N}$ can be thought of as a set of i.i.d. samples drawn from the joint distribution $p(\vec{\theta},\vec{\lambda})=p(\vec{\theta}|\vec{\lambda})p_{\lambda}(\vec{\lambda})$. Under these considerations, the expected KL divergence can be shown to equal the following sum over training datapoints (up to an additive constant):
@@
which can be used for tractable density evaluation, data generation, and augmentation. In other words, the trained emulator can, in principle, facilitate SBI and feature prediction for state-of-the-art simulators such as POSYDONv2, both of which are demonstrated in figure 1, for a BBH PopSynth dataset that uses earlier versions of POSYDON to simulate common-envelope evolution of isolated stellar binaries (S. S. Bavera et al., 2020).
Figure 1: Density estimation (left) for SBI, and data augmentation (right) for feature-prediction with a MAF-based PopSynth emulators. For the right pannel we focus on under-represented regions of parameter space. Each curve represents the effective spin distribution of BBHs in the test set that belong to an unequal mass ratio and high chirp mass sub-population.
While S. Colloms et al. (2025) have extensively discussed SBI using MLE flows, we show here another astrophysical application of PopSynth emulation. Say one is interested in predicting effective spin distributions of a sub-population of BBHs formed through CE, astrophysical parameters same as our test set, that have unequal mass ratio (<0.4<0.4) and high chirp mass ($7M_{\odot}$). The CE channel struggles to populate this region of the parameter space (M. Zevin et al., 2020). If as few as $O(10^{4})$ binaries are simulated for this grid point, only $O(10)$ systems fall in this region of parameter space, which are not sufficient for accurate feature prediction.
On the other hand, an NF trained on multiple population runs, each with $10^{4}$ binaries corresponding to different grid points can be used to accurately amplify the dataset of interest, even if it is excluded from training, without having to run a larger PopSynth simulation. In Fig. 1, the mentioned spin distribution is shown for the small simulation set, the amplified set obtained using an NF trained on all available simulations except the one being amplified, and a much larger PopSynth rerun ($10^{6}$ binaries) which is treated as the ground truth. It can be seen that the emulator predicts the distribution with reasonable accuracy despite being trained on a collection of highly downsampled populations. In other words, NFs can be used to predict population features of rarely observed systems without having to simulate a very large number of binaries, which can be prohibitive. Similar advantages are expected to arise in the context of feature prediction with POSYDONv2 for CBCs formed from isolated stellar binaries with high metallicities or high natal kicks (K. Belczynski et al., 2010, 2016).
@@ 5.1 Uncertainties in flow predictions
The KL divergence as a function of $\vec{\omega}$ can be complicated with multiple peaks, troughs, and saddle points, leading to gradient-based minimization algorithms finding slightly different answers for the best-fit estimator every time the model is re-trained. Loss function morphology can also change significantly between alternate realizations of thinned and noisy training data. In other words, if we train the same model on the same dataset twice, or on different realizations of a noisy dataset, we won’t necessarily get the exact same answer for the predicted population. It is important to quantify and marginalize over this variance in flow predictions for accurate SBI and feature prediction, particularly in the cases of sparse and noisy training datasets and sub-optimal architecture choices. Hence, instead of a best-fit emulator that deterministically yields a single distribution, it is desirable to predict a range of density functions that are likely to contain the true population with a certain posterior probability given training data, model assumptions, and a priori expectation on DNN parameters. Before we show how that can be achieved, it is worthwhile to discuss the types and sources of these uncertainties that we wish to quantify in a bit more detail.
@@
On the other hand, uncertainty can also arise from the inherent randomness, sparsity, and class-confusion present in the dataset, which is irreducible even with the addition of more training samples (W. He et al., 2025; V. Nemani et al., 2023; E. Hüllermeier & W. Waegeman, 2021). Key examples of these aleatoric/data uncertainties in the PopSynth of merging compact binaries are the sparsity of certain regions of parameter space, such as extreme-mass ratio binaries with “mass-gap” compact objects (M. Zevin et al., 2020; R. Abbott et al., 2020) or CBCs emerging from high metallicity stellar binaries, where the merger rate is expected to be low, degenerate population features shared by different values of $\vec{\lambda}$, and the inherent randomness resulting from numerical interpolation schemes used by simulators like POSYDONv2. They would become particularly relevant when POSYDONv2 is used for SBI from catalogs including mass-gap NSBH-like systems, of which two have already been observed (R. Abbott et al., 2020; A. G. Abac et al., 2024), and for making tractable theoretical predictions for similar or other rarely synthesizable sub-populations. Quantifying data uncertainties in deterministic DNN predictors can be achieved through Bayesian inference (J. Arbel et al., 2023) and parametric or non-parametric modeling of the distribution of noise in the data, as well as the use of deep probabilistic models such as secondary NFs and generative adversarial networks (E. Hüllermeier & W. Waegeman, 2021; J. Whang et al., 2021; S. Dirmeier et al., 2023). Again, such approaches have not been widely used to quantify the uncertainties of NF parameters, which is relevant in scenarios where the NF itself is the underlying DNN predictor.
In the following, we present a new implementation of the Bayesian NF, a generative probabilistic model constructed from BNNs, which can self-consistently quantify epistemic uncertainties that arise from inefficient training and aleatoric uncertainties that arise from the sparsity and stochasticity inherent to the dataset. No approximations for the log-likelihood function are utilized, and rigorous sampling methods are used for Bayesian inference. The resulting predictions for an emulated PopSynth simulation are hence marginalized over such uncertainties and also automatically enable Bayesian comparison among different NF models, thereby minimizing the epistemic uncertainties that arise from choosing sub-optimal architectures and families of networks.
6 Uncertainty Quantification and Model Comparison: Bayesian Normalizing Flows
To quantify these epistemic and aleatoric uncertainties in the NF model for thinned datasets, we propose the following. As can be seen from Eqn. (3), The KL divergence is proportional to the negative log-likelihood of the model parameters given the training dataset (up to an additive constant):
@@ 6 Uncertainty Quantification and Model Comparison: Bayesian Normalizing Flows
where $\log\hat{p}(\vec{\theta},\vec{\lambda}|\vec{\omega})=\log\hat{p}(\vec{\theta}|\vec{\lambda},\vec{\omega})+\log p_{\lambda}(\vec{\lambda})$. Note that this is the exact log-likelihood of the parameters characterizing a probabilistic (NF) model, and no a priori assumption about the distribution of noise in the dataset is necessary to construct its functional form. The training method in Sec. 5 above, therefore, amounts to finding a maximum likelihood density estimate (MLE). To quantify uncertainties, we instead propose to sample the posterior distribution of the model parameters using some suitably chosen prior and the exact log-likelihood:
@@
where m stands for marginalized. This amounts to constructing and training a normalizing flow whose transformations comprise a BNN. Alternatively, the predicted population can be represented as a credible interval instead of a single distribution, which is expected to contain the true population with a certain posterior probability, given training data and prior expectations on the model parameters. The direct application of this framework to SBI and data augmentation is outlined in Sec. 7.
@@ 6.1 Prior choices, Implementation and Scalability
Choosing the prior distribution on model parameters correctly is crucial for convergent and scalable Bayesian inference, particularly for models of such high complexity. To fully explore the likelihood function, highly restrictive priors should be avoided. However, broad priors such as standard Gaussians or broad uniform distributions might lead to slow convergence for highly complex probabilistic models such as discrete NFs that are usually characterized by $O(10^{5}-10^{6})$ parameters. For highly complex models such as MAFs, it is instead desirable to choose a prior that is only supportive near the MLE parameters and yet broad enough for likelihood-driven inference. Furthermore, instead of fixing the width of the prior, it can also be allowed to vary within a certain range. We chose the following simple choice for the prior distribution for this proof-of-concept demonstration.
@@
We start by choosing a maximal scale $\vec{\sigma}=\{\sigma_{0}\}_{i=1}^{n_{\omega}}$, which amounts to assuming a priori that the highest likelihood model parameters are within a factor of at most $\sigma_{0}$ of $\vec{\omega}_{0}$. we then define the standardized model parameters $\vec{\omega}_{st}\underset{i.i.d}{\sim}\mathcal{U}(-1,1)$ which we assume to be distributed uniformly within (1,1)(-1,1). we then rescale these to become numbers that lie within a factor of $\vec{\sigma}$ of $\vec{\omega}_{0}$ and hence construct the prior on $\vec{\omega}=\vec{\omega}_{0}(1+\vec{\sigma}\vec{\omega}_{st})$ as an affine transformed uniform distribution. The prior scale $\vec{\sigma}$ can, in principle, be sampled as well from a uniform or half-normal distribution, but for simplicity, we fix it to a value that allows for a reasonably broad prior and likelihood-driven inference. As we demonstrate shortly, calibration statistics can be constructed for choosing optimal values of $\sigma_{0}$.
@@
While these priors might seem restrictive as compared to a standard Gaussian, they can be expected to suffice as long as the posterior is found to be more tightly constrained, away from the prior boundaries. We show that this is the case in the context of our population synthesis dataset and suggest increasing $\sigma_{0}$ for scenarios wherein the posterior is found to be railing against the prior. More sophisticated prior choices that are less restrictive and yet consistent with fast convergence are left as a future exploration.
@@
Even with these prior choices, the Bayesian NFs can quickly become intractable for high-dimensional distributions since discrete flows like MAF require highly complex networks to accurately capture all the features in the population. Bayesian inference can then amount to exploring (and storing) potentially thousands of copies of nearly a million-parameter model, which is expected to become prohibitive. As we show later on in appendix A.1, CNFs can solve this problem by reducing the required model complexity for accurate density estimation by several orders of magnitude. Since the proposed Bayesian framework is agnostic of model architecture, it can be expected to work with CNFs in the same way as discrete flows, and thereby quantify uncertainties in the emulators of higher-dimensional population distributions. On the other hand, if reducing model complexity through the use of CNFs is not a viable alternative, or if complex embedding nets are necessary to reduce the dimensionality of $\vec{\lambda}$, approximate Bayesian inference can be implemented through a combination of stochastic variational inference and importance sampling which is outlined in appendix A.2.
We leave the full realization of both of these resolutions for the future and restrict to Bayesian MAFs that emulates a two-dimensional synthetic population, and a four-dimensional one, in this proof-of-concept demonstration. As mentioned, the high complexity of the MAF makes Bayesian inference with fully un-restricted priors, and large datasets, infeasible. This also results in BICs that are likely driven more by the change in complexity than that in the maximum likelihood achieved during sampling, regardless of overfitting, given the size of the datasets necessitated by tractable inference. We are, however, able to demonstrate that exact posterior sampling can still facilitate likelihood-driven inference that is more informative than the prior, well-calibrated UQ, and BICs that penalize overtly complex models.
6.2 Calibration of Uncertainties
@@
To appraise the accuracy of the estimated uncertainties and optimize the choice of $\sigma_{0}$, we construct metrics that compare the empirical coverage of a test set to the theoretical coverage predicted by Bayesian inference (S. Bieringer et al., 2024). After drawing M posterior samples of flow parameters, we split them up into m random subsets: $S_{\vec{\omega}}=\bigcup_{i=1}^{m}S_{\vec{\omega}}^{i}$, each representative of a smaller Bayesian inference run. For every posterior sample in a particular subset, we generate a million binaries $\vec{\theta}\in S_{\vec{\theta}}^{ij}=\{\vec{\theta}_{k}\sim\hat{p}(\vec{\theta}|\vec{\lambda}_{test},\vec{\omega}_{j})|\vec{\omega}_{j}\in S_{\vec{\omega}}^{i}\}_{k=1}^{10^{6}}$. we then construct $n_{Q}$ equal probability mass $\vec{\theta}$ bins (quantile binning) from the test set and compute the count of predicted samples in each bin. We then compute credible intervals for the counts and estimate the empirical coverage as the fraction of times (out of m and averaged over bins) the credible interval encloses the actual count of test samples. The calibration curve is then constructed as the relationship between the empirical coverage and the nominal coverage (confidence level of the credible intervals, S. Bieringer et al., 2024). Note, however, that for very large values of $\sigma_{0}$, if all of the walkers get stuck at a local maxima of the likelihood function, there will be biases resulting in low coverage. In other words, low coverage can be indicative of both under-estimation of uncertainties due to small $\sigma_{0}$ or biases due to very large $\sigma_{0}$, causing difficulty in convergence. This issue was not discussed in S. Bieringer et al. (2024).
For comparison, we also compute calibration curves for an ensemble of flows. We re-train emulators using the MLE method on alternate realizations of downsampled datasets (of the same size as the ones used to train the Bayesian flow) and compute the empirical coverage of the test set with respect to the credible intervals yielded by the ensemble. We show that even for ensembles comprised of hundreds of emulators, the Bayesian method is better calibrated than with regard to the true distribution.
@@
[…]
@@
20:  return PPif $\frac{\vec{\omega}-\vec{\omega}_{0}}{\vec{\sigma}\times\vec{\omega}_{0}(\textsc{elementwise})}\in[-1,1]$else -$\infty$
@@
23:  return $\textsc{LogLikelihood}(\vec{\omega},\mathrm{True})$
@@
7 Application: Simulation-based inference and data augmentation with UQ
As shown by S. Colloms et al. (2025), an NF emulator can be used to construct the likelihood of the simulation initial conditions given data $\vec{d}_{o}$ from an observed population of systems, using Bayesian hierarchical inference. However, since they train their flows using the MLE method, their inference is always conditional on a fixed value of the flow parameters $\vec{\omega}=\vec{\omega}_{0}$. We instead propose to use the posterior samples of the flow-parameters to obtain empirical constraints on $\vec{\lambda}$ that are marginalized over flow uncertainties using the following posterior distribution:
$p(\vec{\lambda}|\vec{d}_{o},\vec{d}_{train})=\pi(\vec{\lambda})\int d\vec{\omega}p(\vec{d}_{o}|\vec{\lambda},\vec{\omega})p(\vec{\omega}|\vec{d}_{train})$ (9)
where $p(\vec{d}_{o}|\vec{\lambda},\vec{\omega})$ is essentially the LHS of Eq. (1) of S. Colloms et al. (2025), for a particular value of $\vec{\omega}$, (which they choose to be the MLE corresponding to a single training run). To infer $\vec{\lambda}$, we propose to categorically select $\vec{\omega}\in S_{\vec{\omega}}$ for each draw of $\vec{\lambda}\sim\pi(\vec{\lambda})$ explored during sampling the latter’s posterior distribution and use the flow corresponding to the selected $\vec{\omega}$ values for computing the likelihood $p(\vec{d}_{o}|\vec{\lambda},\vec{\omega})$.
Similarly, in the context of data augmentation for rarely synthesizable sub-spopulations, we propose the following. Instead of using the MLE flows trained on a grid of simulations to draw a large number $\vec{\theta}$ samples, we suggest generating an ensemble of augmented datasets, one corresponding to each posterior draw of the flow parameters: $\vec{d}_{A}(\vec{\lambda}_{test})=\{\vec{\theta}_{ij}\sim\hat{p}(\vec{\theta}|\vec{\lambda}_{test},\vec{\omega}_{j})|\vec{\omega}_{j}\in S_{\vec{\omega}}\}$. This dataset can be used to provide a credible interval of the predicted density function for the sub-population in question, which will contain the true prediction as well as that of a single MLE flow with a certain posterior probability. We show this in our demonstration.